OpenStudy (anonymous):

Let f(x) = 5x + 12. Find f–1(x).

OpenStudy (anonymous):

y=5x+12 y-12=5x x=1/5(y-12) x=f-1(y) f-1(y)=1/5(y-12) f–1(x).=1/5(x-12)